OpenStudy (anonymous):

A 12-oz cup of regular coffee at Starbucks has 13 times the amount of caffeine found in the same sized serving of decaffeinated coffee. Together they contain 280 mg of caffeine. How much caffeine is in each type of coffee? Keep trying to figure this out but I keep getting it wrong....

OpenStudy (anonymous):

x is reg y is decaf x+y=280 13y=x 13y+y=280 14y=280 y=20 x=260 check 20*13=260

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

oh okay I see what I did wrong. Thanks!

OpenStudy (anonymous):

No problem:D